# H3AE: High Compression, High Speed, and High Quality AutoEncoder for Video Diffusion Models

### v1

@@ Abstract
Autoencoder (AE) is the key to the success of latent diffusion models for image and video generation, reducing the denoising resolution and improving efficiency. However, the power of AE has long been underexplored in terms of network design, compression ratio, and training strategy. In this work, we systematically examine the architecture design choices and optimize the computation distribution to obtain a series of efficient and high-compression video AEs that can decode in real time on mobile devices. We also unify the design of plain Autoencoder and image-conditioned I2V VAE, achieving multifunctionality in a single network. In addition, we find that the widely adopted discriminative losses, i.e., GAN, LPIPS, and DWT losses, provide no significant improvements when training AEs at scale. We propose a novel latent consistency loss that does not require complicated discriminator design or hyperparameter tuning, but provides stable improvements in reconstruction quality. Our AE achieves an ultra-high compression ratio and real-time decoding speed on mobile while outperforming prior art in terms of reconstruction metrics by a large margin. We finally validate our AE by training a DiT on its latent space and demonstrate fast, high-quality text-to-video generation capability.
**footnotetext: Equal contribution.
@@ 1 Introduction
Over the past year, diffusion models have revolutionized content creation by significantly improving generation quality and expanding text-to-image (T2I) generation [47, 45, 12, 4, 3, 2] to the more complex text-to-video (T2V) task [6, 71, 46, 40]. Recent advances from both industry [46, 59, 6] and the research community [65, 71, 55] now enable the creation of cinematic videos from text prompts [40, 59] or image inputs [5, 46]. These models also empower broader applications such as video editing [25, 36] and 3D generation [60, 33].
@@
Figure 1: Compression ratio is in $log_{2}$ scale. Our H3AE not only achieves better compression ratio vs. PSNR tradeoff, but is also faster and parameter efficient. Refer Tab. 5 for more benchmarks.
The main powerhouse behind this success is Latent Diffusion Modeling (LDM) [47, 39]. Unlike pixel diffusion models [41, 50, 21, 10] which learns the diffusion mapping from noise to raw pixel space, LDMs learn the diffusion process in a compressed latent representation provided by a Variational Autoencoder (VAE) [30, 66, 1], which results in significantly better training and inference efficiency. Quite a few work have been done to improve the generation efficiency and mobile deployment of LDMs, which either compresses the denoising network [34, 23, 70, 62], or reduces the denoising steps [32, 64, 68]. However, the potential of VAEs has long been underexplored from the efficiency perspective. Intuitively, for a transformer denoiser, reducing the size of the latent space by a factor of $O(n)$ would result in theoretical gains of up to $O(n^{2})$ in computational cost during inference. This speed up is of utmost importance to video generation tasks considering the extra temporal dimension. Furthermore, as mentioned in [34, 62], VAE decoding emerges as a significant speed bottleneck in the generation process after successfully optimizing the denoiser by reducing its computation and steps, posing the demand for making VAE efficient while preserving reconstruction quality.
Previous work [9, 1, 18] has explored various aspects of VAE design in an ad-hoc manner, without a comprehensive study on architecture choices and training strategies. Key open questions include: (i) Compression: VAE compression significantly boosts the speed of diffusion models, especially for video generation. What is the highest achievable VAE compression ratio and how can we obtain better Pareto optimality for compression vs. quality? (ii) Efficiency: Given that VAE operates at pixel space, and thus is at risk of running slow or overflowing memory, how to design the network architecture to best balance performance and efficiency? (iii) Quality: How to effectively train video VAEs, i.e., what kind of augmentations and auxiliary losses can enhance their reconstruction quality? In this work, we provide a comprehensive study to address these questions.
Optimize for High Compression & High Speed VAE. We perform an ablation study on design choices (normalization, upsampling, latent channels, etc.) to decide the best VAE architecture. We find that 3D Causal Convolution in the entire network, especially in the high resolution stages, is neither efficient nor necessary for reconstruction quality. Instead, we use 2D convolutions to form disentangled spatial stages to balance quality and efficiency. We show that this is empirically better than using a non-parameterized patchify layer as in LTX [18]. Decoupling spatial stages allows us to slice and perform frame-wise encoding and decoding at high resolution stages. We further introduce 3D Causal Attention in video Autoencoder to achieve better performance with even less parameters. With the optimized design space, we perform mobile latency profiling and distribute the computation accordingly to form an efficient VAE backbone.
Training Algorithm to Improve Quality. Our proposed novel training algorithm consists of the following two components:
@@
  • (a) Omni-Objective. Recently, Reducio [57] proposed a VAE design for image-to-video generation by leaking the first frame to the decoder to improve reconstruction quality with a higher compression ratio compared to T2V. To leverage this advantage and still support a single VAE for T2V and I2V generations, we propose a novel design that unifies decoding irrespective of the presence of image condition. We randomly drop the hierarchical features of the first frame as a condition to the decoder during training, such that our VAE can optionally act as an I2V VAE with improved quality. Surprisingly, this training strategy improves reconstruction even without image conditioning (plain T2V setting) due to the effect of training augmentation.
  • (b) Latent Consistency Loss. Existing works [65, 31, 18, 61] mainly use L1 reconstruction loss, KL regularization, LPIPS loss, and GAN loss to train the VAE. We show that these discriminative losses provide little performance improvements but significantly slow down training. We propose a new training strategy where we first train only with reconstruction loss and KL regularization for faster convergence. Then, we utilize a novel latent consistency loss to further finetune the model, by re-encoding the reconstructed video to obtain the fake posterior and optimize its KL divergence against the former posterior encoded from the ground truth video. This simple loss does not require complicated discriminator design but improves video reconstruction fidelity, both qualitatively and quantitatively.
@@
  • We systematically investigate the design of video VAEs, and propose an efficient architecture that achieves high compression ratio and high inference speed, while maintaining high reconstruction quality.
  • We propose to inject and randomly dropout the first frame condition to train a single VAE supporting both T2V and I2V tasks. The proposed method not only gives us a stronger I2V VAE, but also boosts the reconstruction quality in the pure T2V setting.
  • We revisit VAE training, proposing a novel, simple, and stable latent consistency loss to boost the reconstruction quality, outperforming all prior losses such as LPIPS, GAN and DWT.
  • We validate our VAE by training a 2B DiT in its highly compressed latent space, and demonstrate the first ever DiT-based video generation on mobile device.
@@ 2 Related Work
Latent Diffusion Models. Diffusion models [58, 47, 52, 20, 43, 29] have emerged as the dominant framework for generative modeling, surpassing traditional approaches like GANs [15, 28] and VAEs [30]. These models learn to iteratively denoise an initial noise sample into semantically meaningful and rich content through user defined input controls such as text-prompt and image. To learn large-scale diffusion models, early attempts such as text-to-image (T2I) [50, 10] and text-to-video(T2V) [41, 21] involved directly mapping the noise to the raw pixel space. While these works demonstrated the potential of diffusion models in generative modeling, running very deep neural networks in raw pixel space is computationally very expensive. This led to the development of Latent Diffusion Models (LDMs) [47, 5], which first learn a bi-directional mapping to compress the raw pixel space into a compressed and semantically rich latent space, followed by learning a diffusion model in this latent space. This helps in reducing the computational cost.
Many state-of-the-art T2I [45, 12, 4, 27, 14, 56, 38, 34, 23] and T2V [6, 71, 41, 46, 18, 65] models follow this latent diffusion modeling paradigm to synthesize visually stunning and semantically aligned content. For instance, SDXL [45] trains a large UNet architecture in the latent space to learn a foundational T2I model. SD3 [12] and Flux [4] improves upon this model by training a purely transformer model in a much more advanced latent space and changing the diffusion process to rectified flow for improved generation quality and faster sampling speed. Similarly, many large foundational T2V models utilize LDM framework for video generation. These include Stable Video Diffusion [5], Sora [6], Mochi [55], CogVideoX [22, 65], MovieGen [46] etc. In SnapGen-V [62], the compression factor of the latent models and reduced denoiser foot-print helps to enable their deployment on mobile devices.
Video Autoencoders. Autoencoders play a crucial role in determining both the quality and efficiency of LDMs, leading the research community to dedicate significant effort to improving their reconstruction performance and compression rates. Early foundational video LDMs [5, 17] were recycling image-based $8{\times}8$-compression AEs for the video domain. Such design was leaving a lot of temporal redundancy in the latent representations, and, to support longer video generation, later works (e.g., [65, 71, 72, 31]) pursued $4{\times}8{\times}8$ compression with the causal temporal structure [66]. Most recent foundational video LDMs explore AEs with higher compression rates of $8{\times}8{\times}$ 8 × 8 [46, 1, 54] or even $8{\times}16{\times}16$ [40]. While modern video AEs follow a typical convolutional design [65, 49, 69], several works explored wavelet transforms [16] to efficiently reduce the expensive representation dimensionality of high resolutions [37, 1, 35]. Other works develop transformer-based architectures for video auto-encoding [13, 9, 19, 67, 8].
@@
3 Method
Spatial Upsample $|z|$ Norm Attention Params (M) FPS@512×512@$\times 512$ 512 PSNR SSIM rFVD
@@ 3 Method
3.1 VAE Architecture
3.1.1 Micro Design
Inspired by most recent video VAEs [65, 1, 18], we start from a plain autoencoder backbone built with 3D Causal Convolutions, and explore the following design choices to construct a powerful yet efficient architecture.
Disentangled Spatial Stages.  Applying Conv3D at every stage of a video autoencoder can result in excessive memory consumption, especially in high-resolution stages, thereby hindering its efficient mobile deployment. To mitigate this issue, we disentangle the autoencoder structure by using 2D spatial convolutions in high-resolution stages, as illustrated in Fig. 2. This disentanglement reduces computation and peak memory for high-resolution stages and, in addition, allows chunked inference by slicing video frames. Notably, disentangling the high-resolution stage with Conv2D only has a negligible impact on the reconstruction quality compared to Conv3D, as shown in Tab. 1, while the non-parameterized patchify introduced by LTX [18] demonstrates degraded performance.
Upsampling. Besides traditional interpolation methods, PixelShuffle has become a popular trend in Autoencoders [9]. It requires additional learnable parameters to expand the feature channels, and performs channel-to-space/time operations as upsampling. We observe that PixelShuffle gives much better reconstruction quality than interpolation, with negligible overhead in parameters and latency, as in Tab. 1.
Latent Channels.  Prior continuous variational autoencoders typically adopt a small number of latent channels, such as $4ch$ in stable diffusion [48]. However, recent work have shown that using more latent channels not only improves reconstruction quality, but also offers better semantic completeness and aesthetic quality for diffusion models [12, 22, 23]. We explore larger number of latent channels for our VAE (i.e. 128-channel and 256-channel) in Tab. 1. We find that both 128-channel and 256-channel give reasonable reconstruction performance, and the 256-channel one achieves relatively higher PSNR. We report video diffusion results for both settings.
Normalization Layers.  Though static BatchNorm is foldable during the inference phase thus most efficient, we find that it fails to generalize to different resolutions for VAE reconstruction. We explore dynamic norms including PixelNorm [28] (PN), LayerNorm (LN) and GroupNorm (GN) in this work. Among them, LN is the default choice in transformer-based models, while GN is popular in CNN-based generative models [22, 65]. We find that PN, LN, and GN have similar reconstruction performance, as in Tab. 1. Note that PN is the simplest without learnable parameters, and in addition, LN and GN require reshape operations to process 5D video data, which are less efficient on mobile devices. As a result, we choose PN as our normalization method.
3.1.2 3D Causal Attention
With the operator selection optimized, we introduce 3D Causal Attention as an alternative foundation block in our VAE. 3D Causal Attention has a global spatial receptive field, while only attends to present and previous frames in temporal dimension. Tokens are flattened from $\mathbb{R}^{T\times H\times W\times C}$ to $\mathbb{R}^{(T\cdot H\cdot W)\times C}$ before being processed by the attention mechanism. To enforce causality, a block causal mask is applied to the attention score, where each block has the size of $L_{H}\times L_{W}$ indicating that the tokens are within a single frame, as illustrated in Fig. 2 (right). The mask ensures that video tokens from future frames remain inaccessible to tokens from earlier frames, which preserves temporal consistency and enables arbitrary-length video processing. In addition, QK-norm and RoPE [53] are also incorporated to improve training stability and convergence speed. Notably, RoPE also enhances the generalization capability to various resolutions, which is crucial for a foundational video Autoencoder. As in Tab. 1, using Causal 3D Attention reduces $38\%$ parameters, but achieves 0.9 higher PSNR.
3.1.3 Efficient and High Compression Autoencoder

### v2

@@ Abstract
Autoencoder (AE) is the key to the success of latent diffusion models for image and video generation, reducing the denoising resolution and improving efficiency. However, the power of AE has long been underexplored in terms of network design, compression ratio, and training strategy. In this work, we systematically examine the architecture design choices and optimize the computation distribution to obtain a series of efficient and high-compression video AEs that can decode in real time even on mobile devices. We also propose an omni-training objective to unify the design of plain Autoencoder and image-conditioned I2V VAE, achieving multifunctionality in a single VAE network but with enhanced quality. In addition, we propose a novel latent consistency loss that provides stable improvements in reconstruction quality. Latent consistency loss outperforms prior auxiliary losses including LPIPS, GAN and DWT in terms of both quality improvements and simplicity. H3AE achieves ultra-high compression ratios and real-time decoding speed on GPU and mobile, and outperforms prior arts in terms of reconstruction metrics by a large margin. We finally validate our AE by training a DiT on its latent space and demonstrate fast, high-quality text-to-video generation capability.
@@ 1 Introduction
Figure 1: Compression ratio is in $log_{2}$ scale. H3AE achieves a better compression–PSNR trade-off and is faster and more parameter-efficient. Refer Tab. 3 for more benchmarks.
Over the past year, diffusion-based generative models have achieved remarkable progress, extending the success of text-to-image (T2I) models (Rombach et al., 2022a; Podell et al., 2023; Esser et al., 2024; Black Forest Labs, 2023; Betker et al., 2023; Baldridge et al., 2024) to the more complex text-to-video (T2V) generation (Brooks et al., 2024; Zheng et al., 2024; Polyak et al., 2024; Ma et al., 2025; Wan et al., 2025). Recent advances from both industry (Polyak et al., 2024; Veo-Team et al., 2024; Brooks et al., 2024) and academia (Yang et al., 2024; Zheng et al., 2024; Team, 2024a) enable the creation of cinematic videos from text prompts (Ma et al., 2025; Veo-Team et al., 2024) or image inputs (Blattmann et al., 2023; Polyak et al., 2024). These foundation models also unlock new applications such as video editing (Jeong et al., 2024; Liang et al., 2023) and 3D generation (Voleti et al., 2024; Kwak et al., 2024).
The main powerhouse behind this success is Latent Diffusion Modeling (LDM) (Rombach et al., 2022a; Liu et al., 2022). Unlike pixel diffusion models (Menapace et al., 2024; Saharia et al., 2022; Ho et al., 2022; DeepFloyd, 2023) which learn the diffusion mapping from noise to raw pixel space, LDMs learn the diffusion process in a compressed latent representation provided by a Variational Autoencoder (VAE) (Kingma & Welling, 2013; Yu et al., 2023; Agarwal et al., 2025), yielding dramatic improvements in both training and inference efficiency. The effectiveness of VAE critically depends on three main factors: (i) Reconstruction quality, imperfect reconstruction introduces artifacts into the generation pipeline. (ii) Compression ratio, which directly determines the efficiency of the denoising process: compressing latent resolution by a factor of $O(n)$ yields up to $O(n^{2})$ computational savings for transformer-based denoisers. (iii) Architectural efficiency, especially for decoding, which has emerged as a speed bottleneck after recent advances in reducing denoiser complexity and denoising steps (Li et al., 2023; Hu et al., 2024; Zhao et al., 2024b; Wu et al., 2024; Kulikov et al., 2023; Xu et al., 2024; Zhang et al., 2024). Despite their centrality, VAEs in diffusion pipelines have received less systematic study compared to denoisers. Existing video VAEs (Agarwal et al., 2025; HaCohen et al., 2024) explored design choices in an ad hoc manner, such as adding input pachifications or bringing/discarding auxiliary losses, but leave open key questions about the compression-quality trade-off, the architectural balance between speed and fidelity, and effective training strategies. In this work, we introduce H3AE, a holistic framework that addresses the two principal domains for video VAE optimization:
From the efficiency perspective, we explore VAE designs in terms of: (a) Decoding Efficiency (VAE architecture): We design a compact, low-latency decoder tailored for video autoencoding. Rather than adopting causal 3D convolutions throughout, we structure the decoder in disentangled stages using 2D Conv, 3D Conv and causal attentions, promoting efficiency and enabling sliced decoding for high resolution stages. Through profiling and ablation, we distribute computation across layers to balance speed, memory, and reconstruction fidelity, resulting in a decoding backbone that is fast enough for real-time decoding or even mobile usage while maintaining strong reconstruction. (b) Denoising Efficiency (Compression Ratio): Thanks to our efficient but powerful architecture, we push the compression ratio further than prior VAEs while still preserving high-quality reconstructions, as in Fig. 1. This more aggressive compression reduces the token count fed into the diffusion model, thereby accelerating denoising and lowering memory costs. To validate that our VAE still supports effective diffusion, we train a video DiT model on its latent space and demonstrate high-quality, fast video generation—evidence of “diffusability”.
@@
On the training algorithm side, we propose two methods to improve quality: (a) Omni-Objective Training:Recently, Reducio (Tian et al., 2024) proposed a VAE design for image-to-video generation by leaking the first frame to the decoder to improve reconstruction quality with a higher compression ratio compared to T2V. To leverage this advantage and still support a single VAE for T2V and I2V generations, we propose a novel design that unifies decoding irrespective of the presence of image condition. We randomly feed the hierarchical features of the first frame as a condition to the decoder during training, such that our VAE can optionally act as an I2V VAE with improved quality. Surprisingly, this training strategy improves reconstruction even without image conditioning (plain T2V setting) due to the effect of training augmentation. (b) Latent Consistency Loss:Existing work (Yang et al., 2024; Kong et al., 2024; HaCohen et al., 2024; Wang et al., 2025) mainly use L1 reconstruction loss, KL regularization, and auxiliary losses (LPIPS, GAN, DWT) to train the VAE. We empirically show that the quality gain from these pre-defined auxiliary losses are limited, and they are at risk of introducing checkerboard artifacts. We propose a new training strategy where we first train only with reconstruction loss and KL regularization for faster convergence. Then, we utilize a novel latent consistency loss to further finetune the model by re-encoding the reconstructed video to obtain the fake posterior and optimize its KL divergence against the former posterior encoded from the ground truth video. In our experiments, latent consistency loss is simple to integrate, stable under training, and consistently improves reconstruction fidelity, outperforming previous auxiliary losses.
@@
  • We systematically explore the design space of video VAEs (normalization, upsampling, temporal vs spatial factorization, causal attention, etc.) and propose a new architecture that realizes strong reconstruction, high compression, and low-latency decoding.
  • We propose an omni-objective training method by randomly injecting first-frame hierarchical features to the decoder during training, enabling one VAE to support both T2V and I2V settings, and improving reconstruction even in the unconditioned T2V setting due to training-time augmentation.
  • We propose a novel latent consistency loss to further boost the reconstruction quality, outperforming prior auxiliary losses such as LPIPS, GAN and DWT. Latent consistency loss is stable and simple to integrate into VAE training.
  • We validate H3AE by training a Video DiT in its highly compressed latent space, demonstrating nice diffusability and fast generation speed.
@@ 2 Related Work
Latent Diffusion Models. Diffusion models (Vahdat et al., 2021; Rombach et al., 2022a; Song et al., 2021; Ho et al., 2020; Nichol & Dhariwal, 2021; Karras et al., 2022) have emerged as the leading framework for generative modeling, surpassing earlier approaches such as GANs (Goodfellow et al., 2014; Karras et al., 2019) and VAEs (Kingma & Welling, 2013). These models generate stunning visuals by progressively denoising noise into meaningful context, guided by inputs such as text prompts or images. Early text-to-image (T2I) (Saharia et al., 2022; DeepFloyd, 2023) and text-to-video (T2V) (Menapace et al., 2024; Ho et al., 2022) diffusion models operated directly in pixel space, but this was computationally prohibitive due to the need for very deep networks. Latent Diffusion Models (LDMs) (Rombach et al., 2022a; Blattmann et al., 2023) addressed this issue by first compressing pixels into a semantically rich latent space via an autoencoder, and then applying the diffusion process in that space. This dramatically reduced computational costs while preserving generative quality. Today, most state-of-the-art image (Podell et al., 2023; Esser et al., 2024; Black Forest Labs, 2023; Kag et al., 2024; Gao et al., 2024; Team, 2024b; Liu et al., 2024; Li et al., 2023; Hu et al., 2024) and video (Blattmann et al., 2023; Brooks et al., 2024; Zheng et al., 2024; Menapace et al., 2024; Polyak et al., 2024; Team, 2024a; HaCohen et al., 2024; Yang et al., 2024; Wu et al., 2024) models adopt this paradigm, achieving visually compelling and semantically aligned outputs.
Video Autoencoders. Since autoencoders define the latent space of LDMs, their design directly impacts both quality and efficiency. Early video LDMs (Blattmann et al., 2023; Guo et al., 2023) reused image-based $8{\times}8$ spatial autoencoders, which failed to compress temporal redundancy. To enable longer videos, later works introduced spatio-temporal compression schemes, such as $4{\times}8{\times}8$ (Yang et al., 2024; Zheng et al., 2024; Zhou et al., 2024; Kong et al., 2024), and causal temporal structures (Yu et al., 2023). More recent models push compression further, exploring $8{\times}8{\times}8$ (Polyak et al., 2024; Agarwal et al., 2025; Tang et al., 2024), $8{\times}16{\times}16$ (Ma et al., 2025), or even $8{\times}32{\times}32$ (HaCohen et al., 2024). While extreme compression unlocks faster diffusion (Xie et al., 2024; Chen et al., 2025b; HaCohen et al., 2024), it risks poor reconstructions and requires careful validation during diffusion training (Skorokhodov et al., 2025).
Architecture. Most video autoencoders rely on convolutional backbones (Yang et al., 2024; Sadat et al., 2024; Zhao et al., 2024a), but alternative designs are emerging. Wavelet-based methods (Graps, 1995; Lin et al., 2024; Agarwal et al., 2025; Li et al., 2024) offer efficient handling of high-resolution data, while transformer-based approaches (Esteves et al., 2024; Chen et al., 2025b; Hansen-Estruch et al., 2025; Yu et al., 2024; Chen et al., 2025a) leverage tokenization for scalable latent representations. Recent 1D image tokenizers (Yu et al., 2024; Chen et al., 2025a) push compression to the extreme (up to $2048{\times}$) by adopting ViT-like (Dosovitskiy et al., 2020) autoencoding backbones, hinting at new possibilities for video representation learning.
3 Method
3.1 VAE Architecture
3.1.1 Micro Design
Inspired by recent video VAEs (Yang et al., 2024; Agarwal et al., 2025; HaCohen et al., 2024), we start from a plain autoencoder backbone built with 3D Causal Convolutions, and explore the following design choices to construct a powerful yet efficient architecture.
Spatial Stage.  Applying Conv3D at every stage of a video autoencoder can result in excessive memory consumption, especially in high-resolution stages, thereby hindering its efficient mobile deployment. To mitigate this issue, we disentangle the autoencoder structure by using 2D spatial convolutions in high-resolution stages, as illustrated in Fig. 2. This disentanglement reduces computation and peak memory for high-resolution stages and, in addition, allows frame-by-frame chunked inference. Notably, disentangling the high-resolution stage with Conv2D only has a negligible impact on the reconstruction quality compared to Conv3D, as shown in Tab. 1, while the non-parameterized patchify introduced by LTX (HaCohen et al., 2024) demonstrates degraded performance.
3D Causal Attention.  We introduce 3D Causal Attention as an alternative foundation block in our VAE. 3D Causal Attention has a global spatial receptive field, while only attends to present and previous frames in the temporal dimension. Tokens are flattened from $\mathbb{R}^{T\times H\times W\times C}$ to $\mathbb{R}^{(T\cdot H\cdot W)\times C}$ before being processed by the attention mechanism. To enforce causality, a block causal mask is applied to the attention score, where each block has the size of $L_{H}\times L_{W}$, indicating that the tokens are within a single frame, as illustrated in Fig. 2 (right). The mask ensures that video tokens from future frames remain inaccessible to tokens from earlier frames, which preserves temporal consistency and enables arbitrary-length video processing. In addition, QK-norm and RoPE are also incorporated to improve training stability and convergence speed. Notably, RoPE also enhances the generalization capability to various resolutions, which is crucial for a foundational video Autoencoder. As in Tab. 1, using Causal 3D Attention reduces $38\%$ parameters, but achieves 0.9 higher PSNR.
Upsampling. Besides traditional interpolation methods, PixelShuffle has become a popular trend in Autoencoders (Chen et al., 2025b). We observe that PixelShuffle gives much better reconstruction quality than interpolation, with negligible overhead in parameters and latency, as in Tab. 1.
Latent Channels.  Prior continuous variational autoencoders typically adopt a small number of latent channels, such as $4ch$ in stable diffusion (Rombach et al., 2022b). However, recent work have shown that using more latent channels not only improves reconstruction quality, but also offers better semantic completeness and aesthetic quality for diffusion models (Esser et al., 2024; Hong et al., 2022; Hu et al., 2024). We explore larger number of latent channels for our VAE (i.e. 128-channel and 256-channel) in Tab. 1. We find that both 128-channel and 256-channel give reasonable reconstruction performance, and the 256-channel one achieves relatively higher PSNR. We report video diffusion results for both settings.
Normalization Layers.  Though static BatchNorm is foldable during the inference phase thus most efficient, we find that it fails to generalize to different resolutions for VAE reconstruction. We explore dynamic norms including PixelNorm (Karras et al., 2019) (PN), LayerNorm (LN) and GroupNorm (GN) in this work. Among them, LN is the default choice in transformer-based models, while GN is popular in CNN-based generative models (Hong et al., 2022; Yang et al., 2024). We find that PN, LN, and GN have similar reconstruction performance, as in Tab. 1. Note that PN is the simplest without learnable parameters, and in addition, LN and GN require reshape operations to process 5D video data, which are less efficient on mobile devices. As a result, we choose PN as our normalization method.
@@ 3.1.1 Micro Design
3.1.2 Macro Architecture